OpenStudy (anonymous):

Solve for x: -5<-3(x-1.5)

OpenStudy (anonymous):

\[-5\leq -3(x-1.5)\leq 2\] \[-5\leq -3x+4.5\leq 2\] \[-9.5\leq -3x\leq -2.5\] lets use fractions \[-\frac{19}{2}\leq -3x\leq -\frac{5}{2}\] divide by -3, change the inequality \[\frac{19}{6}\geq x \geq \frac{5}{6}\]

OpenStudy (anonymous):

I had it correct until the fraction part. I had it in decimal form. Thanks